# Balance Correlations, Agentic Zeros, and Networks: The Structure of 192 Years of War and Peace

### v1

@@ 3.1 Data
To demonstrate the usefulness and flexibility of this model, we will use the Correlates of War (CoW) data. These data are the result of an ambitious project started in 1969 by a political scientist at the University of Michigan, David Singer (see Singer,, 1988; Bremer et al.,, 2022, for a current and exhaustive review). This data set includes a comprehensive accounting of both formal defense alliances and military acts of aggression that occurred each year between all pairs of countries from 1816 to 2007. We considered any formal defense alliance between the pair of countries as a positive tie. Acts of aggression recorded in the data set included many categories, ranging from threats of military action against another country to all out wars or invasions (see Table 6 taken from Palmeri et al., 2020). We coded any aggressive act by one country against another as a negative tie. In the rare case where both a defense alliance and an act of aggression occurred within the same year, we coded that tie as a negative tie for that year.
As detailed in Palmeri et al., (2020), there were 22 CoW codes regarding negative ties between nations. They clustered these categories into five buckets, each representing a level of hostility between nations, where 1 represented no military action and 5 represented a full scale war. Instances of no military action (level 1 per their coding scheme) were not coded as a negative tie in our analysis. The remaining 21 codes, however, were deemed a negative tie in this paper. See Table 6 for a full listing of these codes.
@@
Threat to use force [2] Show of force [3] Occupation of territory [4] Begin interstate war [5]
Threat to blockade [2] Alert [3] Seizure [4] Join interstate war [5]
Threat to occupy territory [2] Nuclear alert [3] Attack [4]
Threat to declare war [2] Mobilization [3] Clash [4]
Threat to use CBR weapons [2] Fortify border [3] Declaration of war [4]
Threat to join war [2] Border violation [3] Use of CBR weapons [4]
Blockade [4]
Note: The above 21 coding categories used in the CoW data are organized into severity buckets, labeled [2] through [5], where [5] is the most severe type of event (see Palmeri et al., (2020) for details). We left out the category ‘No militarized action [1]’ as this did not reasonably indicate a negative tie.
@@ 3.2 Testing Balance Correlations
Testing a balance correlation raises the question of which null distribution, against which we would compare the observed statistic, is pertinent.
@@
We rely on a Monte Carlo-simulation to generate a sample of random networks from a data generating process (DGP) assumed for the population. This DGP preserves the marginal distribution of negative, positive and zero ties observed in the data for each year. For example, if a country has a defense alliance with 12 other countries, and engages in an act of aggression with 5 other countries, then the DGP would generate sample networks that would preserve this expected number of positive and negative ties with other countries. In addition, our null-generating process preserves the observed covariance in distribution of negative, positive and zero ties in each year. This constitutes a conservative test, one that restricts the sampling space support taking the empirical marginals of the data matrices as expected degree. This degree based approach is more conservative than for example an Erdös-Renyi random network approach, because the latter would allow a wider range of possible random networks to be considered. In our data set of country ties, we assume that there is diversity between countries in their capacity to engage in positive and negative ties. Note that this doesn’t imply countries with less capacity are not making active choices. Their choice sets differ, but we assume they can and do consider all possible ties explicitly or implicitly. This assumption supports our choice to base the test on degree expectations, as it includes the investment distribution all countries make.
A further issue to consider is the fact that the randomly generated matrices are mutually exclusive in their elements, i.e. the sum over the primitive matrices must be exactly one. The procedure we take is that for each element in the network we sample one of three values $\{-1,0,1\}$, with probability based on expected row and column degrees of each of the N, Z, and P networks, respectively. This process converges in the limit to observed degree distribution of the empirical data set.
@@
This DGP approach is based on the observed data for each given year. Thus, the set of possible null graphs can be seriously constrained, and these constraints can change for each balance correlation for each year. The resulting distribution of null graphs may exhibit unfamiliar shapes. For example, in Figure 3 we see the distribution of balance correlations for the set of <pp> 2-path balance correlations (<p.pp>, <n.pp>, and <z.pp>). The <p.pp> balance correlations are mostly positive, ranging from just below zero to just over .5. The <z.pp> balance correlations, however, are mostly negative, ranging from -.4 to about .1. The distribution for the <n.pp> balance correlations, however, is much more peaked, with a pronounced mode of about -.03 but a range from just below that up to +.18. The shape of each of these distributions depends on the number of countries in the data set that year, the density of each type of tie, the degree distributions of each type of tie, and the covariances among these degree distributions. These peculiarities are the reason that an appropriate DGP was used to test each balance correlation separately for each year.
Figure 3: Example of Distribution of Balance Correlations under the Null
@@
Balance Correlation Note: In this example, one thousand null graphs were generated from a model that simulated networks that had expected p, n, and z degrees equal to those observed in 1916. The observed densities of the positive, negative and zero ties were .023, .036, and .940, respectively. This histogram displays the distribution of three balance correlations (<p.pp>, <n.pp>, and <z.pp>) calculated from these null graphs. The blue-shaded part of the histogam, mostly on the right side of the chart, displays the frequency distribution of the <p.pp> correlations. The yellow-shaded part, mostly on the left side of the chart, displays the frequency distribution of the <z.pp> correlations. The red-shaded portion of the histogram, the more peaked distribution in the middle of the chart, displays the frequency distribution of the <n.pp> correlations. The solid vertical colored lines represent the corresponding balance correlations for the observed data for that year. The observed balance correlations for both <p.pp> (the vertical blue line) and <z.pp> (the vertical yellow line) were outside the null region, suggesting they were statistically significant coefficients. The observed balance correlation <n.pp> (the vertical red line), however, was within the null distribution region and therefore not statistically significant.
@@ 3.3 Results
Note: The three dotted plots show how the density of the positive, negative and zero ties varied over the 192 years studied. The blue dots indicate the density of positive ties, the red dots indicate the density of negative ties, and the black dots indicate the density of the zero ties. These densities within any given year by definition sum to 1.0. Of special note is the period of time between 1867 and 1936, the LOA years, during which there was a precipitous decline in the density of positive ties (alliances) between nations.
In Figure 4, we plot the density of ties of each of the three types over this time period. First and foremost, it is clear that Z-ties are the most prevalent, with densities ranging from .75 to .99 over these 192 years. Negative ties are usually (but not always) less frequent than positive ties, with spikes in density most notably during WWI and WWII. Perhaps most interesting is the pattern of positive ties. In general, the density of international relations characterized by formal defense alliances of some type varied between almost none (1909) to almost 20% (at the conclusion of WWII). What is striking in this plot is that these alliances were most frequent before 1867 and after 1936. In between these two years, the density of inter-nation defense agreements dropped considerably. 1867 was the beginning of the dissolution of the Austro-Hungarian empire; 1936 was the onset of WWII with the re-militarization of the Rhineland. We will call this period the “Lack of Alliances” (or LOA) period. We use it here to refer to this specific span of years (1867 to 1936), one that we found to be critical in our exploration of balance, given the regularity with which patterns of balance were interrupted during this period. It can be characterized as a period not of relative calm as much as a period of disorder, where alliances were either not sought or not lasting.
@@
In each of the following figures, we plot the balance correlations for each year across the 192 years for which we have data (1816 to 2007). In addition, we show the 95% interval from the DGP simulated null distribution in gray. Thus, points representing balance correlations outside that gray region can be considered statistically significant in that they would not be expected to appear under the null model. These significant balance correlations are indicated with a red boundary around the blue dot. Balance correlations that are within the bounds of the null model, thus not deemed statistically significant, are indicated with a small blue dot only.
@@ 3.3.1 PP Family
Figure (5) contains the three plots of the <pp> family of balance correlations for <p.pp>, <n.pp>, and <z.pp>. In the first plot, we examine the pattern of the classic balance claim that a “friend of a friend is a friend.” We see that the <p.pp> balance correlations For the most part are persistently strong, ranging mostly from .8 to .95, providing strong evidence for this balance condition. The primary exception to this lies within the LOA period we identified earlier. During this time, we can see that the country level behavior in the world tends to result in less balance on a global level. It is clear, also, that the low number of positive alliances gives rise to a wider band of balance correlations under the null model. This suggests system wide turmoil.
@@ 3.3.2 PN (and Twin NP) Family
Figure 6, showing the three plots of balance correlations in the PN family, provides an interesting extension of the results in the prior figure. While the balance correlations for <p.pn>, the first panel in Figure 6, show a predominent tendency towards correlations near zero up to the years just prior to WWII, the period subsequent to WWII contain mostly significantly positive correlations. This is true despite the fact that <p.pn> configuration theoretically represents an unbalanced condition: One should not consider an enemy of a friend (or its twin condition, a friend of an enemy) a friend. Yet it seems to be so, and steadily so since the late 1950’s.
On the other hand, the second plot in Figure 6 shows reasonable evidence for balance in that for a majority of years the <n.pn> configuration is also positive and significant. And the strength of these findings, as represented by the size of the correlations, exceeds the size of the balance-disconfirming correlations in the <p.pn> plot. That is, during much of this period of time we examined, the propensity to engage in a hostile act against a country who is a friend of an enemy is substantially greater than if they are not a friend of an enemy. The fact that both of these conditions (<p.pn> and <n.pn>) exist at the same time speaks to the complexity of international relations. To presume that only one balance rule operates in any given temporal or network circumstance is belied by these results.
What is also clear is the z-ties tend to be significantly suppressed in the face of <pn> 2-path relations. Again, it suggests neutrality is less an option many times in the face of these 2-path conditions. This general tendency, though, does not always coincide with the years during which this was true for the <z.pp> balance correlations in Figure 5.
@@ 3.3.3 PZ (and Twin ZP) Family
Figure 7 is our first case where all members of the family are excluded from any of Heider’s original predictions. Thus, it is difficult to say whether any of these correlations should be theoretically balanced. What we do observe, though, is whether they appear to be empirically balanced. That is, are there any general patterns that show persistently positive (balanced) correlations, or perhaps persistently negative (unbalanced) correlations?
@@ 3.3.4 NN Family
Figure 8 provides the balance correlations for the <nn> family. The first plot in this figure provides general support of the Heiderian balance proposition that “an enemy of an enemy is a friend” in that for over half of the years the <p.nn> correlation is positive and significant. In the post LOA period, since WWII, the <z.nn>, the <z.nn> correlation is almost always negative, and often significantly so, suggesting that neutral stances are perhaps discouraged with an enemy of an enemy.
Of interest in the <z.nn> and <n.nn> plots is the large bands areas of gray, which represent the range of correlations one is likely to observe if the null hypothesis were true. In particular, in the <n.nn> plot, the null region often ranges from a correlation of zero up to 1.0. While most of the observed correlations are either 0 or close to it, the wide range of possible null values suggests that we should not be confident in much of anything in this plot. This is largely because the density of n-ties is low to begin with, and thus the density of <nn> 2-path ties to be even less frequent. This scarcity permits the correlation to be quite unstable under the null hypothesis.
@@ 3.3.5 NZ (and Twin ZN) Family
Figure 9 shows the results for the <nz> (or <zn> twin) family. Almost all of these balance correlations are either zero of close to zero, with the minor exception of the significantly positive <n.nz> correlations during the two world wars. Not only are the preponderance of these balance correlations near zero, they also are either just barely border the null region or clearly not significant at all. Thus, we have no evidence that a country who is an enemy of a neutral partner is likely to differentially engage in any particular relation (p, n or z) with another nation.
@@ 3.3.6 ZZ Family
Our final set of results are provided in Figure 10. Z ties, and <zz> 2-path ties, are the most dense. Thus, as is displayed in the plots, the null regions are relatively thin. The <n.zz> balance correlations are still largely insignificant, either inside or bordering the null region.
@@
It is worth deliberating on the meaning of these <zz> structures. These suggest that the Ego  Alter dyad is surrounded by non-ties or neutral ties to an inordinate number of other countries. It is as if Ego and Alter are in separate or separated worlds. Thus, a <p.zz> pattern would suggest that Ego and Alter tend to form a bridge between these relatively disconnected parties. A positive balance correlation for <p.zz> would indicate that an alliance is more likely if they are from these distant parts; a negative correlation would indicate that Ego and Alter are less likely to form an alliance if they are from these separated parts. What we see from the <p.zz> plot is that the balance correlations are generally if only slightly negative, suggesting alliances are discouraged across these chasms.
@@
A <z.zz> triple tells a slightly different story. In this case, a positive correlation would indicate that Ego  Alter pairs across such gulfs are relatively more likely to stay neutral with each other. This would suggest that there are relatively disconnected clusters in the community, and that inside these clusters are where much of the action (both positive and negative) are happening. Indeed, the <z.zz> balance correlations, again outside the LOA period, are positive, if not exceptionally strong.
@@
What is interesting in this family of plots, however, is that for the <p.zz> and <z.zz> balance correlations outside the LOA years, we find corelations that are significant (not contained in the null region) yet closer to zero than the null region itself. This pattern is particularly striking in the post LOA years, during and following WWII. The <p.zz> correlations are predominantly negative and yet they are significantly greater than we would expect by chance under the null.
4 Discussion and Extensions

### v2

@@ 3 Empirical example
3.1 Background
The importance of Balance Theory for international relations has been widely recognized [35, 1, 13, 30, 47]. Current models of (structural) balance theory allow for consideration of either two types of relations (positive, or negative) (see [22, 3]). Any further refinement in strength, or nominal categorization of relationships is impossible within the current framework. The proposed approach in this paper does allow to investigate many more refinements, which occur in international relations. Note however, that although we are studying data over time, this is not a study into the dynamics of international relations. Rather we identify balance correlations per year that subsequently substantiate a time series. In other words we study embedded triadic balance behaviour at a system level, not the feedback of system level changes into the considerations of countries when making international relational choices, although this is a further extension we would highly support, it is not studied in this paper.
@@ 3.2 Data
To demonstrate the usefulness and flexibility of this model, we will use the Correlates of War (CoW) data. These data are the result of an ambitious project started in 1969 by a political scientist at the University of Michigan, David Singer [44, 4] (for a current and exhaustive review). This data set includes a comprehensive accounting of both formal defense alliances and military acts of aggression that occurred each year between all pairs of countries from 1816 to 2007. We considered any formal defense alliance between the pair of countries as a positive tie. Acts of aggression recorded in the data set included many categories, ranging from threats of military action against another country to all out wars or invasions (see Table 6 taken from [40]). We coded any aggressive act by one country against another as a negative tie. In the rare case where both a defense alliance and an act of aggression occurred within the same year, we coded that tie as a negative tie for that year.
As detailed in [40], there were 22 CoW codes regarding negative ties between nations. They clustered these categories into five buckets, each representing a level of hostility between nations, where 1 represented no military action and 5 represented a full scale war. Instances of no military action (level 1 per their coding scheme) were not coded as a negative tie in our analysis. The remaining 21 codes, however, were deemed a negative tie in this paper. See Table 6 for a full listing of these codes.
@@
Code $2$ Code $3$ Code $4$ Code $5$
Threat to use force Show of force Occupation of territory Begin interstate war
Threat to blockade Alert Seizure Join interstate war
Threat to occupy territory Nuclear alert Attack
Threat to declare war Mobilization Clash
Threat to use CBR weapons Fortify border Declaration of war
Threat to join war Border violation Use of CBR weapons
Blockade
Note: The above 21 coding categories used in the CoW data are organized into severity buckets, labeled $2$ through $5$, where $5$ is the most severe type of event (see [40] for details). We left out the category ‘No militarized action’ Code $1$ as this did not reasonably indicate a negative tie.
Under the assumption of limited relational resources it is reasonable to argue that countries will consider relations with the whole set of other countries. They are limited in developing ties due to resources restrictions, which differ per country. It is recognized that the implicit or explicit neutrality choice is an embedded triad choice driven by a set of possible motives including geographical distance, but render this a political decision.
@@ 3.3 Testing balance correlations
Testing balance correlation raises the question: against which null distribution? We rely on a Monte Carlo-simulation to generate a sample of random networks from a data generating process (DGP) assumed for the population. This DGP preserves the marginal distribution of negative, positive and zero ties observed in the data for each year. For example, if a country has a defense alliance with 12 other countries, and engages in an act of aggression with 5 other countries, then the DGP would generate sample networks that would preserve this number of positive and negative ties with other countries in expectation.
To construct such a DGP, we make use of the result that the maximum-entropy model with specific expectations for some functions of the random variable of interest is (to use modern statistical terminology) an exponential family model. Its sufficient statistic must comprise the functions whose expectations are to be fixed [29]. Exponential-family random graph models (ERGMs) have been developed both for scenarios in which positive and negative ties may co-occur and must thus be modeled jointly [28] and for scenarios in which they cannot [19], with the general polytomous relationship case first specified by [43]. Our application then requires the sufficient statistic vector to be the concatenation of positive and negative tie degrees of each node, resulting in a model of the form
$P(\mathcal{G};\boldsymbol{\eta})=\exp({\boldsymbol{\eta}\cdot[\textbf{1}_{n}^{\top}\textbf{P},\textbf{1}_{n}^{\top}\textbf{N}]^{\top}})/{\kappa(\boldsymbol{\eta})},$ (9)
where $\textbf{1}_{n}$ is an $n$-vector of 1s and $\kappa(\boldsymbol{\eta})$ is the normalising constant so that $P(\mathcal{G};\boldsymbol{\eta})$ sums to 1 over the sample space of signed networks of interest. Estimation and simulation are then possible using general techniques for valued ERGMs [33], or by a specialized implementation. In a parallel development, this and other null models for signed graphs were studied by [20], including an optimized implementation.
@@
In addition, this null-generating process all preserves the observed covariance in distribution of negative, positive and zero ties in each year. This constitutes a conservative test, one that restricts the sampling space support taking the empirical marginals of the data matrices as expected degree. This degree based approach is more conservative than for example an Erdös-Renyi random network approach [20], because the latter would allow a wider range of possible random networks to be considered. In our data set of country ties, we assume that there is diversity between countries in their capacity to engage in positive and negative ties. Note that this doesn’t imply countries with less capacity are not making active choices. Their choice sets differ, but we assume they can and do consider all possible ties explicitly or implicitly. This assumption supports our choice to base the test on degree expectations, as it includes the investment distribution all countries make.
@@
This DGP approach is based on the observed data for each given year. Thus, the set of probable null graphs can be seriously constrained, and these constraints can change for each balance correlation for each year. The resulting distribution of null graphs may exhibit unfamiliar shapes. For example, in Fig 3 we see the distribution of balance correlations for the set of <pp> 2-path balance correlations (<p.pp>, <n.pp>, and <z.pp>). The <p.pp> balance correlations are mostly positive, ranging from just below zero to .60.60.60.60. The <z.pp> balance correlations, however, are mostly negative, ranging from .$-.55$ to .00.00.00.00 The distribution for the <n.pp> balance correlations, ranges from .$-.05$ but a range from just below that up to just below +.$+.40$. The shape of each of these distributions depends on the number of countries in the data set that year, the density of each type of tie, the degree distributions of each type of tie, and the covariances among these degree distributions. Especially for the latter two distributions this leads to stron bi-modal distributions as can be seen from Fig 3. These peculiarities may differ per year. Therefore a separate DGP was used to test each balance correlation for each year.
Figure 3: Example of balance correlations under the null distribution
@@
Note: In this example, $1000$ graphs were generated from a null-model that simulated networks that had expected p, n, and z degrees equal to those observed in 1916. The observed densities of the positive, negative and zero ties were .023.023.023.023, .036.036.036.036, and .940.940.940.940, respectively. This histogram displays the distribution of three balance correlations (<p.pp>, <n.pp>, and <z.pp>) calculated from these null graphs. The green-shaded part of the histogram, mostly on the right side of the chart, displays the frequency distribution of the <p.pp> correlations. The teal-shaded part, mostly on the left side of the chart, displays the frequency distribution of the <z.pp> correlations. The purple-shaded portion of the histogram, in the middle of the chart, displays the frequency distribution of the <n.pp> correlations. The dashed vertical colored lines represent the corresponding balance correlations for the observed data for that year. The observed balance correlations for both <p.pp> (the vertical green line) and <z.pp> (the vertical teal line) were outside the null region, suggesting they were statistically significant coefficients. The observed balance correlation <n.pp> (the vertical purple line), however, was within the null distribution region and therefore not statistically significant.
@@ 3.4 Results
Figure 4: Log of number of nodes, dyads, and $p$, $n$ and $z$ ties per year
@@
Note: The three dotted plots show how the density of the positive, negative and zero ties varied over the 192 years studied. The green dots indicate the density of positive ties, the purple dots indicate the density of negative ties, and the teal dots indicate the density of the zero ties. These densities within any given year by definition sum to 1.0. Of special note is the period of time between 1867 and 1936, the LOA years, during which there was a precipitous decline in the density of positive ties (alliances) between nations.
In Fig 5, we plot the density of ties of each of the three types over this time period. First and foremost, it is clear that Z-ties are the most prevalent, with densities ranging from .75 to .99 over these 192 years. Negative ties are usually (but not always) less frequent than positive ties, with spikes in density most notably during WWI and WWII. Perhaps most interesting is the pattern of positive ties. In general, the density of international relations characterized by formal defense alliances of some type varied between almost none (1909) to almost 20% (at the conclusion of WWII). What is striking in this plot is that these alliances were most frequent before 1867 and after 1936. In between these two years, the density of inter-nation defense agreements dropped considerably. 1867 was the beginning of the dissolution of the Austro-Hungarian empire; 1936 was the onset of WWII with the re-militarization of the Rhineland. We will call this period the “Lack of Alliances” (or LOA) period. We use it here to refer to this specific span of years (1867 to 1936), one that we found to be critical in our exploration of balance, given the regularity with which patterns of balance were interrupted during this period. It can be characterized as a period not of relative calm as much as a period of disorder, where alliances were either not sought or not lasting.
@@
In each of the following figures, we plot the balance correlations for each year across the 192 years for which we have data (1816 to 2007). In addition, we show the 95% interval from the DGP simulated null distribution in gray. Thus, points representing balance correlations outside that gray region can be considered statistically significant in that they would not be expected to appear under the null model. Balance correlations that are within the bounds of the null model, thus are not deemed statistically significant.
@@ 3.4.1 PP family
Fig (6) contains the three plots of the <pp> family of balance correlations for <p.pp>, <n.pp>, and <z.pp>. In the first plot, we examine the pattern of the classic balance claim that a “friend of a friend is a friend.” We see that the <p.pp> balance correlations For the most part are persistently strong and significant, ranging mostly from .8 to .95, providing strong evidence for this balance condition. The primary exception to this lies within the LOA period we identified earlier. During this time, we can see that the country level behavior in the world tends to result in less balance on a global level, while it can also not be inferred that the observed balance structure is due to balance oriented behavior. It is clear, also, that the low number of positive alliances gives rise to a wider band of balance correlations under the null model. This suggests system wide turmoil.
@@ 3.4.2 PN (and twin NP) family
Fig 7, showing the three plots of balance correlations in the PN family, provides an interesting extension of the results in the prior figure. While the balance correlations for <p.pn>, the first panel in Fig 7, show a predominent tendency towards correlations near zero up to the years just prior to WWII, the period subsequent to WWII contain mostly significantly positive correlations. This is true despite the fact that <p.pn> configuration theoretically represents an unbalanced condition: One should not consider an enemy of a friend (or its twin condition, a friend of an enemy) a friend. Yet it seems to be a consistent strategy since the late 1950’s. Establishing these structures is a tertius gaudens strategy [5] .
On the other hand, the second plot in Fig 7 shows evidence for balance in that for a majority of years the <n.pn> configuration is also positive, significant and never negative over this period. And the strength of these findings, as represented by the size of the correlations, exceeds the size of the balance-disconfirming correlations in the <p.pn> plot. That is, during much of this period of time we examined, the propensity to engage in a hostile act against a country who is a friend of an enemy is substantially greater than if they are not a friend of an enemy. The fact that both of these conditions (<p.pn> and <n.pn>) exist at the same time speaks to the complexity of international relations. To presume that only one balance rule operates in any given temporal or network circumstance is belied by these results.
What is also clear is the z-ties tend to be significantly suppressed in the face of <pn> 2-path relations. Again, it suggests neutrality is less an option many times in the face of these 2-path conditions. This general tendency, though, does not always coincide with the years during which this was true for the <z.pp> balance correlations in Fig 6.
@@ 3.4.3 PZ (and twin ZP) family
Fig 8 is our first case where all members of the family are excluded from any of Heider’s original predictions. Thus, it is difficult to say whether any of these correlations should be theoretically balanced. What we do observe, though, is whether they appear to be empirically balanced. That is, are there any general patterns that show persistently positive (balanced) correlations, or perhaps persistently negative (unbalanced) correlations?
@@ 3.4.4 NN family
Fig 9 provides the balance correlations for the <nn> family. The first plot in this figure provides general support of the Heiderian balance proposition that “an enemy of an enemy is a friend” in that for over half of the years the <p.nn> correlation is positive and significant. In the post LOA period, since WWII, the <z.nn> correlation is almost always significantly negative, suggesting that neutral stances are perhaps discouraged with an enemy of an enemy.
Of interest in the <n.nn> plots is the large range of correlations resulting from the null model. In particular, in the <n.nn> plot, the null region often ranges from a correlation of zero up to .95. While most of the observed correlations are either 0 or close to it. This is largely because the density of n-ties is low to begin with, and thus the density of <nn> 2-path ties to be even less frequent. This scarcity leaves the correlation to be quite unstable under the null hypothesis.
@@ 3.4.5 NZ (and twin ZN) family
Fig 10 shows the results for the <nz> (or <zn> twin) family. Almost all of these balance correlations are either zero of close to zero, with the minor exception of the significantly positive <n.nz> correlations during the two world wars. Not only are the preponderance of these balance correlations near zero, they also are either just barely border the null region or clearly not significant at all. Thus, we have no evidence that a country who is an enemy of a neutral partner is likely to differentially engage in any particular relation (p, n or z) with another nation.
@@ 3.4.6 ZZ family
Our final set of results are provided in Fig 11. Z ties, and <zz> 2-path ties, are the most dense. Thus, as is displayed in the plots, the null regions are relatively thin. The <n.zz> balance correlations are still largely insignificant, either inside or bordering the null region.
@@
It is worth deliberating on the meaning of these <zz> structures. These suggest that the Ego  Alter dyad is surrounded by non-ties or neutral ties to an inordinate number of other countries. It is as if Ego and Alter are in separate or separated worlds. Thus, a <p.zz> pattern would suggest that Ego and Alter tend to form a bridge between these relatively disconnected parties. A positive balance correlation for <p.zz> would indicate that an alliance is more likely if they are from these distant parts; a negative correlation would indicate that Ego and Alter are less likely to form an alliance if they are from these separated parts. What we see from the <p.zz> plot is that the balance correlations are generally if only slightly negative, suggesting alliances are discouraged across these chasms, but occur more frequently than expected under the null-model.
@@
A <z.zz> triple tells a slightly different story. In this case, a positive correlation would indicate that Ego  Alter pairs across such gulfs are relatively more likely to stay neutral with each other. This would suggest that there are relatively disconnected clusters in the community, and that inside these clusters are where much of the action (both positive and negative) are happening. Indeed, the <z.zz> balance correlations, again outside the LOA period, are positive, if not exceptionally strong, but occur less frequently than expected. This supports the idea of a push towards global integration.
@@
What is interesting in this family of plots, however, is that for the <p.zz> and <z.zz> balance correlations outside the LOA years, we find correlations that are significant (not contained in the null region) yet closer to zero than the null region itself. This pattern is particularly striking in the post LOA years, during and following WWII. The <p.zz> correlations are predominantly negative and yet they are significantly greater than we would expect by chance under the null, while the opposite holds for the <z.zz> correlations.
Figure 12: Empirical vs theoretical Heiderian balance
3.4.7 Overview of results
Fig 12 provides a visual summary of the results across all years. Each point in the plot represents the average balance correlation across all 192 years for which there are data. They are divided into three groups: the “theoretically balanced" group (<p.pp>, <n.pn>, and <p.nn>), the “theoretically unbalanced" group (<n.pp>, <p.pn>, and <n.nn>), and the “indeterminate" group composed of the remaining 12 triad signature configurations that contain at least one z-tie.
Several overall findings are apparent from this figure. First, the strongest finding supports the original Heiderian balance prediction represented by the <p.pp> signature. That is, at this international level, countries who have alliances with common third countries tend to have alliances with each other: “ A friend of a friend is a friend." The other theoretically predicted balanced structures are weaker, but at least some of the <n.pn> structures are significant over time (see Fig 7). Interestingly, none of Heider’s theoretically predicted unbalanced structures receive empirical support. All of these correlations hover around zero. Indeed, if anything, the <n.nn> and <p.pn> unbalanced structures show relatively frequent positive correlations, just the opposite of what Heiderian theory would predict.
Perhaps the most interesting, and surprising, results involve those configurations with a z-tie. The most striking negative correlation does not emerge from any of the theoretically unbalanced triples but rather occurs in the <z.pp> configuration. That is, being an ally of an ally seems to carry no systematic information about hostile acts between two countries that differentiates it from other contexts. In contrast, a neutral tie is far less likely to occur in this context than in any other context. Other z-tie results also engage in unexpected patterns, although the strength of those findings is not as evident as is the <z.pp> case. We dilate these results and based on the correlations among Balance Correlations make the case that beside the theoretical arguments for treating z-ties as separate category, the empirical results strongly corroborate this idea.
3.4.8 Interpreting z-ties
In our example on relations between countries, clusters of balance theoretical behaviors can be identified, i.e. clusters of balance correlations can be discerned that are positively correlated, irrespective of whether the balance correlations are significant or not. Furthermore, these clusters may show negative correlations with balance correlations in other clusters. Earlier we discussed the marking difference between <p.pp> and <z.pp>. Fig 13 shows a strong negative correlation (.$-.95$) between these two signatures. So the mean values of these balance correlation are opposite, and this is consistent over time. An even stronger negative correlation occurs between <z.pz> and <p.pz> (.$-.99$). Although the balance correlations themselves are not extreme, the observed signatures in our example seem to be highly dependent. It is stressed that this is an empirical effect not a priori determined.
Using a hierarchical clustering, looking at $7$ clusters gives a first singleton <n.nn> (unbalanced), and $4$ pairs (<n.pn> and <p.nn>, both balanced; <n.pp> and <p.pn>, both unbalanced; and two indeterminate pairs, <n.nz> and <p.zz>, and, <z.pn> and <z.nn>). The balanced and unbalanced pairs have much stronger correlations, .86.86.86.86 and .79.79.79.79, respectively, compared to .31.31.31.31 and .48.48.48.48 for the respective indeterminate pairs. There are two larger blocks with 4 and 5 signatures, respectively. The first contains signatures from 4 different families, and includes the balanced <p.pp>, which is most strongly correlated with <z.pz> (.75.75.75.75), while the other two, <z.nz> and <n.zz> are even stronger correlated (.91.91.91.91) the other correlations range from .24.24.24.24 to .39.39.39.39. In the largest cluster all signatures are theoretically indeterminate with highest correlations between the pair <n.pz> and <p.nz> (.93.93.93.93), and the pair <z.pp> and <p.pz> (.78.78.78.78). The other correlations in this cluster range from .23.23.23.23 and .53.53.53.53.
Now this result may helpful to assess whether the z-ties should also empirically be considered a separate type. To interpret the valence of z-ties as either positive or negative if we replace these z’s with either a p or n designation, such that it provides a balance theoretical prediction consistent with the observed mean balance correlation (Fig  12).
Figure 13: Correlations of balance correlations over 192 years
Table 7: Establishing ambiguity of z-ties.
Indeterminate signature $\overline{BC}$ (Un)Balanced imputed signature Correlation with the imputed signature (CWIS) z-tie ambiguity?
<z.pn> .$-.12$ <p.pn> .$-.61$ Yes, CWIS
<z.nn> .$-.10$ <n.nn> .$-.45$ Yes, CWIS
<z.nz> .01.01.01.01 <n.pn> or <p.nn> .$-.24$ or .$-.22$ Yes, imput. ambig.
<n.zz> .$-.03$ <n.nn> or <n.pp> Both not sig. Yes, CWIS
<z.pz> .10.10.10.10 <n.pn> or <p.pp> Not sig. or .75.75.75.75 No, z could be p
<z.pp> .$-.72$ <n.pp> Not sig. Yes, CWIS
<p.pz> .$-.10$ <p.pn> .$-.35$ Yes, CWIS
<z.zz> .12.12.12.12 <p.pp> Not sig. Yes, CWIS
<n.pz> .$-.01$ <n.pp> .32.32.32.32 No, z could be p
<p.nz> .$-.03$ <p.pn> Not sig. Yes, CWIS
<n.nz> .05.05.05.05 <n.pn> Not sig. Yes, CWIS
<p.zz> .$-.11$ <p.pn> .32.32.32.32 Yes, imput. ambig.
Note: Bold printed are $z$-imputations. In the first $z$-imputation of $n$ and $p$ in <z.nz>, instead of <n.np> the correlation with the twin signature <n.pn> is reported. In imputing $p$ in <p.nz> instead of <p.np> the correlation with twin signature <p.pn> is reported. In imputing $p$ in <n.nz> instead of <n.np> the correlation with twin signature <n.pn> is reported.
In Table 7 little support for the idea that z-ties can be considered either as positive or negative is found. Rather there is much support for what we call z-tie ambiguity. When z-labels are imputed with n or p labels, according to achieve a signature that is coherent with Heiderian balance theory multiple situations can be distinguished. First, in most cases coherence can be achieved by exclusive p or n imputation. Second, coherence can be brought about by either p or n imputation. Third, coherence can only be achieved by imputation of both p and n.
[…]
In the second case where a coherent mean balance correlation occurs whenever either an n or an p gets imputed there is a case for which no ambiguity can be detected. In the case of <z.pz> when z is imputed with the p label, a strong positive correlation (.75.75.75.75) between <z.pz> and <p.pp> supports the conjecture that z-ties are to be seen as positive ties. However, in the other indeterminate signature such support is not found as correlation of <n.zz> with both <n.pp> as well as with <n.nn> are not significant.
[…]
4 Extensions